OpenStudy (anonymous):

A 50-foot tree casts a shadow 75 feet long. How can you find sine for the angle of elevation that is determined by the sun and the top of the tree?

OpenStudy (anonymous):

c2=502+752 c2=5625+2500 c2=8125 c=90.14 sin(θ)=50/90.14 sin(θ)=0.55